k = ab + (c/2)

find the value of k when a = 3 , b = 4, c = 6
1
Expert's answer
2012-10-08T07:59:51-0400
k = ab + (c/2)
a = 3 , b = 4, c = 6 : k=3*4+6/2=12+3=15